# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 2.1 Datalog
where $\psi_{0}$ is an atomic $\rho$-formula and $\psi_{1},\dots,\psi_{n}$ are atomic $(\rho\cup\tau)$-formulas; we also assume that all rules are safe, i.e., that every variable that appears in the head also appears in the body. If $\mathfrak{A}$ is a $\tau$-structure, and $\Pi$ is a Datalog program with EDBs from $\tau$ and IDBs $\rho$, then a $(\tau\cup\rho)$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ is called a fixed point of $\Pi$ on $\mathfrak{A}$ if $\mathfrak{A}^{\prime}$ satisfies the sentence
@@
for each rule $\psi_{0}\;{:}{-}\;\psi_{1},\dots,\psi_{n}$. If $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$ are two $(\rho\cup\tau)$-structures with the same domain $A$, then $\mathfrak{A}_{1}\cap\mathfrak{A}_{2}$ denotes the $(\rho\cup\tau)$-structure with domain $A$ such that
@@
There is an equivalent ‘operational’ definition of the semantics of Datalog, and in particular of $\Pi(\mathfrak{A})$; roughly speaking, we compute $R^{\Pi(\mathfrak{A})}$ for $R\in\rho$ ‘bottom up’, starting from the empty relation, and adding tuples that must be contained in all fixed points, until we reach a (the) smallest fixed point (for details see, e.g., Libkin’s book [38, Section 10.5]; an explicit treatment of the equivalence can e.g. be found in [8, Theorem 8.1.6]). From this equivalent description of $\Pi(\mathfrak{A})$ it is apparent that if $\mathfrak{A}$ is a finite structure then $\Pi(\mathfrak{A})$ can be computed in polynomial time in the size of $\mathfrak{A}$.
If $R\in\rho$, we also say that $\Pi$ defines $R^{\Pi(\mathfrak{A})}$ on $\mathfrak{A}$. A Datalog program together with a distinguished predicate $R\in\rho$ may also be viewed as a formula, which we also call a Datalog query, and which over a given $\tau$-structure $\mathfrak{A}$ denotes the relation $R^{\Pi(\mathfrak{A})}$. If the distinguished predicate has arity 00, we often call it the goal predicate; we say that $\Pi$ derives ${\sf goal}$ on $\mathfrak{A}$ if ${\sf goal}^{\Pi(\mathfrak{A})}=\{()\}$. The class ${\mathcal{C}}$ of finite $\tau$-structures $\mathfrak{A}$ such that $\Pi$ derives ${\sf goal}$ on $\mathfrak{A}$ is called the class of finite $\tau$-structures defined by $\Pi$, and denoted by $\llbracket\Pi\rrbracket$. Note that this class ${\mathcal{C}}$ is definable in universal second-order logic (we have to express that in every expansion of the input by relations for the IDBs that satisfies all the rules of the Datalog program the goal predicate is non-empty).
2.2 Second-Order Logic
@@
For IDBs $X$ and $Y$ we write $Y\subsetneq X$ as a shortcut for
@@
Note that if $\mathfrak{A}$ is an $\{E,D\}$-structure and $X\subseteq A$ satisfies $\psi_{X}$, then any homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ must be constant on $X$. Also note that if $X,Y\subseteq A$ satisfy $\psi_{X}$ and $\psi_{Y}$ and $X\cap Y\neq\emptyset$, then $Z:=X\cap Y$ satisfies $\psi_{Z}$. Hence, for every element $a\in A$ there exists a unique smallest set $X=X_{a}$ that satisfies $\psi_{X}$ and contains $a$.
@@
Indeed, suppose that $\mathfrak{A}$ is an $\{E,D\}$-structure and let $X\subseteq A$ be smallest with the property that it satisfies $\phi_{X}$. If there are $x,y\in X$ such that $(x,y)\in D^{\mathfrak{A}}$ then clearly there is no homomorphism from $\mathfrak{A}$ to $({\mathbb{N}};E,D)$. On the other hand, if $\mathfrak{A}$ satisfies (1), then any map $h\colon A\to{\mathbb{N}}$ such that $h^{-1}(h(a))=X_{a}$ for all $x\in A$ is a homomorphism from $\mathfrak{A}$ to $({\mathbb{N}};E,D)$. Clearly, there exists such a map $h$, which proves the claim.
@@ 2.3 Guarded Second-Order Logic
There are many semantically equivalent ways of introducing GSO [33]. Let $\mathfrak{B}$ be a $\tau$-structure. Then $(t_{1},\dots,t_{n})\in B^{n}$ is called guarded in $\mathfrak{B}$ if there exists an atomic $\tau$-formula $\phi$ and $b_{1},\dots,b_{k}$ such that $\mathfrak{B}\models\phi(b_{1},\dots,b_{k})$ and $\{t_{1},\dots,t_{n}\}\subseteq\{b_{1},\dots,b_{k}\}$. Note that (for $n=1$) every element of $B$ is guarded (because of the atomic formula $x=x$). A relation $R\subseteq B^{n}$ is called guarded if all tuples in $R$ are guarded. Note that all unary relations are guarded. If $\Psi$ is an arbitrary second-order sentence, we say that a structure $\mathfrak{B}$ satisfies $\Psi$ with guarded semantics, in symbols $\mathfrak{B}\models_{g}\Psi$, if all second-order quantifiers in $\Psi$ are evaluated over guarded relations only. Note that for MSO sentences, the usual semantics and the guarded semantics coincide.
@@
It follows that GSO is at least as expressive as MSO. There are many Datalog programs that are equivalent to a GSO sentence, but not to an MSO sentence. However, since MSO is surprisingly expressive (see Example 2.2) it can be quite challenging to prove that a specific problem is inexpressible in MSO. E.g., it is an open problem whether $\operatorname{CSP}({\mathbb{Q}};B)$ is expressible in MSO, where
@@
is the so-called betweenness relation [24]. The proof of the following proposition is based on a variant of an example of a Datalog query in GSO given in [18] (Example 2).
@@
On the left of Figure 1 one can find an example of a $\{S,T,R,N\}$-structure $\mathfrak{B}$ where the given Datalog program derives ${\sf goal}$.
[…]
@@
To show that ${\mathcal{C}}$ is not MSO definable, suppose for contradiction that there exists an MSO sentence $\Phi$ such that $\llbracket\Phi\rrbracket={\mathcal{C}}$. We use $\Phi$ to construct an MSO sentence $\Psi$ which holds on a finite word $w\in\{a,b\}^{*}$ (represented as a structure with signature $P_{a},P_{b},<$ in the usual way [38]) if and only if $w\in\{a^{n}b^{n}\mid n\geq 1\}$; this contradicts the theorem of Büchi-Elgot-Trakhtenbrot (see, e.g., [38]). Let $\Phi^{\prime}$ be the MSO sentence obtained from $\Phi$ by replacing all subformulas of $\Phi$ of the form
  • $S(x,y)$ by a formula $\phi_{S}(x,y)$ that states that $x$ is the smallest element with respect to $<$, that $P_{b}(y)$, and that there is no $z<y$ in $P_{b}$;
  • $T(x,y)$ by a formula $\phi_{T}(x,y)$ that states that $P_{a}(x)$, that there is no $z>x$ in $P_{a}$, and that $y$ is the largest element with respect to $<$;
@@
  • $N(x,y)$ by a formula $\phi_{N}(x,y)$ stating that $y$ is the next element after $x$ w.r.t. $<$.
The resulting MSO sentence $\Psi_{1}$ has the signature $\{P_{a},P_{b},<\}$; let $\Psi$ be the conjunction of $\Psi_{1}$ with the sentence $\Psi_{2}$ which states that for all $x,y\in A$, if $x<y$ and $P_{a}(y)$ then $P_{a}(x)$.
@@
Let $\mathfrak{B}$ be the $\{S,T,R,N\}$-structure such that for $X\in\{S,T,R,N\}$ we have $X^{\mathfrak{B}}:=\{(x,y)\mid\mathfrak{A}\models\phi_{X}(x,y)\}$. See Figure 1 for an example of a structure $\mathfrak{A}$ such that $w_{\mathfrak{A}}=a^{4}b^{4}$ and the corresponding $\{S,T,R,N\}$-structure $\mathfrak{B}$. If $w_{\mathfrak{A}}$ is of the form $a^{n}b^{n}$ for some $n\geq 1$, then $\mathfrak{A}$ clearly satisfies $\Psi_{2}$. To show that it also satisfies $\Psi_{1}$, let $v_{1},\dots,v_{n},w_{1},\dots,w_{n}\in A$ be such that $\{v_{1},\dots,v_{n}\}=P_{a}^{\mathfrak{A}}$ and $\{w_{1},\dots,w_{n}\}=P_{b}^{\mathfrak{A}}$ such that for all $i,j\in\{1,\dots,n\}$, if $i<j$ then $v_{i}<^{\mathfrak{A}}v_{j}$ and $w_{i}<^{\mathfrak{A}}w_{j}$. Then
@@
[…]
It follows that $\mathfrak{B}$ satisfies $\Phi$ and therefore $\mathfrak{A}\models\Psi$.
For the converse direction, suppose that $\mathfrak{A}\models\Psi$. Clearly, $w_{\mathfrak{A}}\in a^{*}b^{*}$ because $\mathfrak{A}\models\Psi_{2}$. Moreover, since $\mathfrak{A}\models\Psi_{1}$ we have that $\mathfrak{B}\models\Phi$, and hence there exist $n\in{\mathbb{N}}$ and elements $v_{1},\dots,v_{n},w_{1},\dots,w_{n}\in A$ such that $\mathfrak{B}$ satisfies (2). We first prove that $P_{a}^{\mathfrak{A}}=\{v_{1},\dots,v_{n}\}$ and $|P_{a}^{\mathfrak{A}}|=n$. Since $(v_{n},w_{n})\in T^{\mathfrak{B}}$ we have $\phi_{T}(v_{n},w_{n})$ and hence $v_{n}\in P_{a}^{\mathfrak{A}}$. Since $\mathfrak{B}\models N(v_{1},v_{2}),\dots,N(v_{n-1},v_{n})$ we have that $v_{1}<v_{2}<\cdots<v_{n-1}<v_{n}$ holds in $\mathfrak{A}$ and it also follows that $|P_{a}^{\mathfrak{A}}|=n$. Then for every $i\in n$ we have that $v_{i}\in P_{a}^{\mathfrak{A}}$ because $v_{i}\leq v_{n}$, $v_{n}\in P_{a}^{\mathfrak{A}}$, and $w_{\mathfrak{A}}\in a^{*}b^{*}$. Now suppose for contradiction that there exists $x\in P^{\mathfrak{A}}_{a}\setminus\{v_{1},\dots,v_{n}\}$; choose $x$ largest with respect to $<^{\mathfrak{A}}$. Since $(v_{n},w_{n})\in T^{\mathfrak{B}}$ and $x\in P_{a}^{\mathfrak{A}}$ we must have $x\leq v_{n}$, and hence $x<v_{n}$ since $x\notin\{v_{1},\dots,v_{n}\}$. Then there exists $y\in A$ such that $\phi_{N}(x,y)$ holds in $\mathfrak{A}$. Since $y\leq v_{n}$, $v_{n}\in P_{a}^{\mathfrak{A}}$, and $w_{\mathfrak{A}}\in a^{*}b^{*}$, we must have $P_{a}^{\mathfrak{A}}$. By the maximal choice of $x$ we get that $y=v_{i}$ for some $i\in\{1,\dots,n\}$. But then $\phi_{N}(x,v_{i})$ implies that $x\in\{v_{1},\dots,v_{n-1}\}$, a contradiction. Similarly, one can prove that $P^{\mathfrak{A}}_{b}=\{w_{1},\dots,w_{n}\}$ and that $|P^{\mathfrak{A}}_{b}|=n$. This implies that $w_{\mathfrak{A}}=a^{n}b^{n}$.
We finally have to prove that ${\mathcal{C}}$ is in GSO. Let $\Phi$ be the GSO $\{S,T,R,N\}$ sentence with existentially quantified unary relations $V,W$, and existentially quntified binary relations $R^{\prime}\subseteq R$ and $N^{\prime}\subseteq N$, which states that

### v3

@@ 2.1 Datalog
where $\psi_{0}$ (called the head of the rule) is an atomic $\rho$-formula and $\psi_{1},\dots,\psi_{n}$ (jointly referred to as the body of the rule) are atomic $(\rho\cup\tau)$-formulas; we also assume that all rules are safe, i.e., that every variable that appears in the head also appears in the body. If $\mathfrak{A}$ is a $\tau$-structure, and $\Pi$ is a Datalog program with EDBs from $\tau$ and IDBs $\rho$, then a $(\tau\cup\rho)$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ is called a fixed point of $\Pi$ on $\mathfrak{A}$ if $\mathfrak{A}^{\prime}$ satisfies the sentence
@@
for each rule $\psi_{0}\;{:}{-}\;\psi_{1},\dots,\psi_{n}$. If $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$ are two $(\rho\cup\tau)$-structures with the same domain $A$, we let $\mathfrak{A}_{1}\cap\mathfrak{A}_{2}$ denote the $(\rho\cup\tau)$-structure with domain $A$ such that
@@
There is an equivalent ‘operational’ definition of the semantics of Datalog, and in particular of $\Pi(\mathfrak{A})$; roughly speaking, we compute $R^{\Pi(\mathfrak{A})}$ for $R\in\rho$ ‘bottom up’, starting from the empty relation, and adding tuples that must be contained in all fixed points, until we reach a (the) smallest fixed point (for details see, e.g., Libkin’s book [40, Section 10.5]; an explicit treatment of the equivalence can e.g. be found in [10, Theorem 8.1.6]). From this equivalent description of $\Pi(\mathfrak{A})$ it is apparent that if $\mathfrak{A}$ is a finite structure then $\Pi(\mathfrak{A})$ can be computed in polynomial time in the size of $\mathfrak{A}$. $\triangle$
If $R\in\rho$, we also say that $\Pi$ defines $R^{\Pi(\mathfrak{A})}$ on $\mathfrak{A}$. A Datalog program together with a distinguished predicate $R\in\rho$ may also be viewed as a formula, which we also call a Datalog query, and which over a given $\tau$-structure $\mathfrak{A}$ denotes the relation $R^{\Pi(\mathfrak{A})}$. If the distinguished predicate has arity 0, we often call it the goal predicate; we say that $\Pi$ derives $\sf goal$} on $\mathfrak{A}$ if ${\sf goal}^{\Pi(\mathfrak{A})}=\{()\}$. The class $\mathcal{C}$} of finite $\tau$-structures $\mathfrak{A}$ such that $\Pi$ derives $\sf goal$} on $\mathfrak{A}$ is called the class of finite $\tau$-structures defined by $\Pi$, and denoted by $\llbracket\Pi\rrbracket$. Note that this class $\mathcal{C}$} is definable in universal second-order logic (we have to express that in every expansion of the input by relations for the IDBs that satisfies all the rules of the Datalog program the goal predicate is non-empty).
2.2 Second-Order Logic and Monadic Second-Order Logic
@@
For unary relation symbols $X$ and $Y$ we write $Y\subsetneq X$ as a shortcut for
@@
Note that if $\mathfrak{A}$ is an $\{E,D\}$-structure and $X\subseteq A$ satisfies $\psi_{X}$, then any homomorphism from $\mathfrak{A}$ to $(\mathbb{N};E,D)$ must be constant on $X$. Also note that if $X,Y\subseteq A$ satisfy $\psi_{X}$ and $\psi_{Y}$ and $X\cap Y\neq\emptyset$, then $Z:=X\cap Y$ satisfies $\psi_{Z}$. Hence, for every element $a\in A$ there exists a unique smallest set $X=X_{a}$ that satisfies $\psi_{X}$ and contains $a$.
@@
Indeed, suppose that $\mathfrak{A}$ is an $\{E,D\}$-structure and let $X\subseteq A$ be smallest with the property that it satisfies $\phi_{X}$. If there are $x,y\in X$ such that $(x,y)\in D^{\mathfrak{A}}$ then clearly there is no homomorphism from $\mathfrak{A}$ to $({\mathbb{N}};E,D)$. On the other hand, if $\mathfrak{A}$ satisfies (1), then any map $h\colon A\to{\mathbb{N}}$ such that $h^{-1}(h(a))=X_{a}$ for all $x\in A$ is a homomorphism from $\mathfrak{A}$ to $({\mathbb{N}};E,D)$. Clearly, there exists such a map $h$, which proves the claim. $\triangle$
@@ 2.3 Guarded Second-Order Logic
There are many semantically equivalent ways of introducing GSO [36]. Let $\mathfrak{B}$ be a $\tau$-structure. Then a tuple $(t_{1},\dots,t_{n})\in B^{n}$ is called guarded in $\mathfrak{B}$ if there exists an atomic $\tau$-formula $\phi$ and $b_{1},\dots,b_{k}$ such that $\mathfrak{B}\models\phi(b_{1},\dots,b_{k})$ and $\{t_{1},\dots,t_{n}\}\subseteq\{b_{1},\dots,b_{k}\}$. A relation $R\subseteq B^{n}$ is called guarded if all tuples in $R$ are guarded. Note that (for $n=1$) every element of $B$ is guarded (because of the valid atomic formula $x=x$), and consequently all unary relations are guarded. If $\Psi$ is an arbitrary second-order sentence, we say that a structure $\mathfrak{B}$ satisfies $\Psi$ with guarded semantics, in symbols $\mathfrak{B}\models_{g}\Psi$, if all second-order quantifiers in $\Psi$ are evaluated over guarded relations only. Note that for MSO sentences, the usual semantics and the guarded semantics coincide.
@@
It follows that GSO is at least as expressive as MSO. There are many Datalog programs that are equivalent to a GSO sentence, but not to an MSO sentence. However, since MSO is surprisingly expressive (see Example 2.2) it may be quite challenging to prove that a specific problem cannot be expressed in MSO. For example, it is an open problem whether $\operatorname{CSP}({\mathbb{Q}};B)$ is expressible in MSO, where
@@
is the so-called betweenness relation [27]. One such MSO-inexpressibility result is part of the following proposition whose proof is based on a variant of an example of a Datalog query in GSO given in [21] (Example 2).
@@
On the left of Figure 1 one can find an example of a $\{S,T,R,N\}$-structure $\mathfrak{B}$ where the given Datalog program derives $\sf goal$}. Note that the Datalog program derives $\sf goal$} on a given finite structure if and only if a ‘ladder’ structure as shown on the left of Figure 1 homomorphically maps to the structure.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$SSTTRRRRNNNNNNNNNNNN
@@
To show that $\mathcal{C}$} is not MSO-expressible, suppose toward a contradiction suppose for contradiction that there exists an MSO sentence $\Phi$ such that $\llbracket\Phi\rrbracket={\mathcal{C}}$. We use $\Phi$ to construct an MSO sentence $\Psi$ which holds on a finite word $w\in\{a,b\}^{*}$ (represented as a structure with the signature $\{P_{a},P_{b},<\}$ in the usual way [40]) if and only if $w\in\{a^{n}b^{n}\mid n\geq 1\}$; this contradicts the theorem of Büchi-Elgot-Trakhtenbrot (see, e.g., [40]), which states that a language of finite words can be expressed in MSO if and only if the language is regular. Let $\Psi_{1}$ be the MSO sentence obtained from $\Phi$ by replacing all subformulas of $\Phi$ of the form
  • $S(x,y)$ by a formula $\phi_{S}(x,y)$ that states that $x$ is the smallest element with respect to <<, that $P_{b}(y)$ holds, and that there is no $z<y$ in $P_{b}$;
  • $T(x,y)$ by a formula $\phi_{T}(x,y)$ that states that $P_{a}(x)$ holds, that there is no $z>x$ in $P_{a}$, and that $y$ is the largest element with respect to <<;
@@
  • $N(x,y)$ by a formula $\phi_{N}(x,y)$ stating that $y$ is the next element after $x$ w.r.t. <<.
The resulting MSO sentence $\Psi_{1}$ has the signature $\{P_{a},P_{b},<\}$. Let $\Psi$ be the conjunction of $\Psi_{1}$ with the sentence $\Psi_{2}$ which states that for all $x,y\in A$, if $x<y$ and $P_{a}(y)$ then $P_{a}(x)$.
@@
Let $\mathfrak{B}$ be the $\{S,T,R,N\}$-structure with the same domain as $\mathfrak{A}$ such that for $X\in\{S,T,R,N\}$ we have $X^{\mathfrak{B}}:=\{(x,y)\mid\mathfrak{A}\models\phi_{X}(x,y)\}$. See Figure 1 for an example of a structure $\mathfrak{A}$ such that $w_{\mathfrak{A}}=a^{4}b^{4}$ and the corresponding $\{S,T,R,N\}$-structure $\mathfrak{B}$. Note that $\mathfrak{B}\models\Phi$ if and only if $\mathfrak{A}\models\Psi_{1}$. If $w_{\mathfrak{A}}$ is of the form $a^{n}b^{n}$ for some $n\geq 1$, then $\mathfrak{A}$ clearly satisfies $\Psi_{2}$. To show that it also satisfies $\Psi_{1}$, let $v_{1},\dots,v_{n},w_{1},\dots,w_{n}\in A$ be such that $\{v_{1},\dots,v_{n}\}=P_{a}^{\mathfrak{A}}$ and $\{w_{1},\dots,w_{n}\}=P_{b}^{\mathfrak{A}}$ such that for all $i,j\in\{1,\dots,n\}$, if $i<j$ then $v_{i}<^{\mathfrak{A}}v_{j}$ and $w_{i}<^{\mathfrak{A}}w_{j}$. Then
@@
$(v_{i},w_{i})\in R^{\mathfrak{B}}$ $\text{ for all }i\in\{2,\dots,n-1\},$ (2)
$(v_{i},v_{i+1}),(w_{i},w_{i+1})\in N^{\mathfrak{B}}$ $\text{ for all }i\in\{1,\dots,n-1\}.$
It follows that $\mathfrak{B}$ satisfies $\Phi$ and therefore $\mathfrak{A}\models\Psi_{1}$.
For the converse direction, suppose that $\mathfrak{A}\models\Psi$. Clearly, $w_{\mathfrak{A}}\in a^{*}b^{*}$ because $\mathfrak{A}\models\Psi_{2}$. Moreover, since $\mathfrak{A}\models\Psi_{1}$ we have that $\mathfrak{B}\models\Phi$, and hence there exist $n\in{\mathbb{N}}$ and elements $v_{1},\dots,v_{n},w_{1},\dots,w_{n}\in A$ such that $\mathfrak{B}$ satisfies (2). We first prove that $P_{a}^{\mathfrak{A}}=\{v_{1},\dots,v_{n}\}$ and $|P_{a}^{\mathfrak{A}}|=n$. Since $(v_{n},w_{n})\in T^{\mathfrak{B}}$ we have $\phi_{T}(v_{n},w_{n})$ and hence $v_{n}\in P_{a}^{\mathfrak{A}}$. Since $\mathfrak{B}\models N(v_{1},v_{2}),\dots,N(v_{n-1},v_{n})$ we have that $v_{1}<v_{2}<\cdots<v_{n-1}<v_{n}$ holds in $\mathfrak{A}$ and it also follows that $|P_{a}^{\mathfrak{A}}|\geq n$. Then for every $i\in n$ we have that $v_{i}\in P_{a}^{\mathfrak{A}}$ because $v_{i}\leq v_{n}$, $v_{n}\in P_{a}^{\mathfrak{A}}$, and $w_{\mathfrak{A}}\in a^{*}b^{*}$. Now suppose for contradiction that there exists $x\in P^{\mathfrak{A}}_{a}\setminus\{v_{1},\dots,v_{n}\}$; choose $x$ largest with respect to $^{\mathfrak{A}}$. Since $(v_{n},w_{n})\in T^{\mathfrak{B}}$ and $x\in P_{a}^{\mathfrak{A}}$ we must have $x\leq v_{n}$, and hence $x<v_{n}$ since $x\notin\{v_{1},\dots,v_{n}\}$. Then there exists $y\in A$ such that $\phi_{N}(x,y)$ holds in $\mathfrak{A}$. Since $y\leq v_{n}$, $v_{n}\in P_{a}^{\mathfrak{A}}$, and $w_{\mathfrak{A}}\in a^{*}b^{*}$, we must have $y\in P_{a}^{\mathfrak{A}}$. By the maximal choice of $x$ we get that $y=v_{i}$ for some $i\in\{1,\dots,n\}$. But then $\phi_{N}(x,v_{i})$ implies that $x\in\{v_{1},\dots,v_{n-1}\}$, a contradiction. Similarly, one can prove that $P^{\mathfrak{A}}_{b}=\{w_{1},\dots,w_{n}\}$ and that $|P^{\mathfrak{A}}_{b}|=n$. This implies that $w_{\mathfrak{A}}=a^{n}b^{n}$.
We finally have to prove that $\mathcal{C}$} is in GSO. Let $\Phi$ be the GSO $\{S,T,R,N\}$ sentence with existentially quantified unary relations $V,W$, and existentially quantified binary relations $R^{\prime}\subseteq R$ and $N^{\prime}\subseteq N$, which states that